# Adaptive Group Robust Ensemble Knowledge Distillation

### v1

@@ 5.2 Results
Teacher $85.7_{\pm 1.80}$ $65.6_{\pm 5.75}$ $95.4_{\pm 0.07}$ $37.5_{\pm 2.27}$ $90.0_{\pm 0.25}$ $75.9_{\pm 1.15}$
$94.2_{\pm 0.53}$ $90.9_{\pm 1.00}$ $93.7_{\pm 2.44}$ $90.1_{\pm 1.68}$ $85.9_{\pm 0.49}$ $77.9_{\pm 0.49}$
Deep Ensemble $84.4_{\pm 0.00}$ $59.3_{\pm 0.44}$ $95.6_{\pm 0.02}$ $37.2_{\pm 0.00}$ $90.6_{\pm 0.02}$ $76.1_{\pm 0.10}$
$93.5_{\pm 0.17}$ $90.0_{\pm 0.56}$ $92.4_{\pm 0.18}$ $90.3_{\pm 0.55}$ $85.9_{\pm 0.17}$ $76.6_{\pm 0.09}$
One-hot $83.8_{\pm 0.97}$ $54.1_{\pm 2.21}$ $95.5_{\pm 0.07}$ $32.3_{\pm 2.66}$ $90.1_{\pm 0.21}$ $75.5_{\pm 0.84}$
$91.4_{\pm{1.31}}$ $86.7_{\pm 1.85}$ $92.3_{\pm 0.39}$ $88.9_{\pm 2.90}$ $85.7_{\pm 0.47}$ $76.5_{\pm 1.09}$
Random $80.0_{\pm 0.37}$ $39.6_{\pm 3.63}$ $95.2_{\pm 0.03}$ $27.2_{\pm 0.00}$ $90.8_{\pm{0.21}}$ $\underline{75.2}_{\pm 0.95}$
@@
AGRE-KD (Ours) $\underline{82.2}_{\pm 1.37}$ $\underline{55.0}_{\pm 5.47}$ $95.4_{\pm 0.04}$ $\underline{37.6}_{\pm 0.78}$ $89.3_{\pm 3.65}$ $74.7_{\pm 3.00}$
$\textbf{91.3}_{\pm 0.49}$ $\textbf{87.9}_{\pm 1.23}$ $91.7_{\pm 0.20}$ $\textbf{91.9}_{\pm 0.71}$ $90.2_{\pm 0.49}$ $\textbf{75.9}_{\pm 1.75}$
We train each model using three random seeds and report the means and standard deviations. We consider ensemble distillation with ten teacher models randomly sampled from a poll of pretrained (biased) teachers across random seeds. In addition to the teacher’s performance and deep ensemble approach, we report the performance of the "student" trained only using the ground truth class labels (One-hot), i.e., without knowledge distillation. We report the overall average and Worst-Group Accuracy (WGA). In the Supplementary C (Table 5), we provide the group-wise accuracy of each baseline method on the Waterbirds and CelebA datasets. We consider models trained using ERM or with last-layer retraining for debiasing. Specifically, for ensemble knowledge distillation methods, we train the student model with biased teachers (trained using ERM) and debiased teachers (last layer retrained with DFR (Kirichenko et al.,, 2022)). Table 1 summarizes the main results of the paper from which we draw the following observations:
  • The deep ensemble using models with the same architecture as the student model improves the worst-case group performance. While for all ensemble knowledge distillation, the worst-case group performance can drop up to 10%. The drop is more severe for the random ensemble distillation but improved average accuracy, which suggests switching between teachers during the training harms worst-case groups. On the other hand, AE-KD (Du et al.,, 2020) performs better than other standard KD methods, showing that addressing disagreement between teachers’ gradient direction can benefit the worst group samples.
  • The fairness property of the last-layer retraining of neural network using DFR  is transferable. When we train the teacher models using ERM and then retrain the classification layer with the group-balanced set, the resulting student models achieve significantly better worst-case group accuracy. We only train the student model using the combined teachers’ outputs without any feature distillation. These results demonstrate that the retrained classifications can provide pseudo-label distribution (dark knowledge) that reduces students’ reliance on spurious features. This shows that by mimicking the teacher’s outputs, the classifier layer of the student model also downweights the spurious features in its last layers (Kirichenko et al.,, 2022). However, the improved students’ WGA across ensemble distillation methods does not match the WGA of the debiased teacher models. We attribute this to the smaller capacity of the student model, which we discuss in the next experiment.
  • Our weighting scheme consistency improves the WGA compared to other ensembling methods. On the CelebA dataset, we achieve better teachers’ average performance in terms of worst-case group accuracy as well as the deep ensemble of models with the same capacity as the student. This observation is consistent with settings where teacher models are biased or debiased. On the other hand, we can see that our method does not provide significant improvement for the worst-case group when all teacher models in the ensemble are biased; we attribute this to the fact all these teachers will likely have a similar gradient direction with biased models for many samples.
Table 2: Results on self-distillation. The student and the teacher models have the same network architecture (resnet50). The student’s worst-group test accuracy increases when we distill to a network with higher capacity.
@@
AGRE-KD $\underline{84.9}_{\pm 1.40}$ $\underline{66.3}_{\pm 4.76}$ $94.5_{\pm 3.22}$ $\underline{39.2}_{\pm 0.78}$
$91.4_{\pm 1.99}$ $\textbf{91.1}_{\pm 2.56}$ $91.1_{\pm 0.09}$ $\textbf{91.9}_{\pm 1.17}$
Impact of the model capacity.
@@
In this experiment, we study whether the student’s network capacity is a source of the implication of spurious feature learning in ensemble KD. In particular, we perform the same experiments as previously but using self-distillation, i.e., we use the same architecture for the student and teacher models (i.e., resnet50); we report the results on Waterbirds and Celeba datasets in Table 2 and provide results with resnet34 student in the Supplementary C (Figure 4). The ensemble KD methods with self-distillation significantly improve the worst-case test group accuracy. The gap between the teacher models and students is reduced compared to KD with a smaller capacity student model (i.e., resnet18 in Table1). The results indicate that the students’ higher capacity can help the network learn more core features and reduce the influence of spurious features in the last layer. On the other hand, AGRE-KD  outperforms other baselines, showing that our adaptive weighting scheme effectively guides the student models to focus on minimizing worst-case group error during training. We further illustrate the effectiveness of our weighting scheme in the next experiment by adjusting the number of debiased teachers in the ensemble.
Effect of the number of debiased teachers in the ensemble.
We study how the WGA of the student model is impacted when the teacher ensemble contains biased and debiased models with different proportions. We consider five teachers in this experiment and use the same training process as previously with different proportions of debiased teachers (i.e., $\{\frac{1}{5}$, $\frac{2}{5}$, …, $\frac{5}{5}\}$); we report the average and WGA in Figure 3 for the Waterbirds and CelebA datasets. The results indicate our method can adaptively identify and rely more on the knowledge of debiased teachers while reducing reliance on biased teachers. AGRE-KD  can significantly improve worst-case performance when the ensemble contains a single debiased teacher. The aggregation process of other ensemble KD methods relies more on biased teachers, leading the student model to capture spurious correlation. As the proportion of debiased teachers increases, the worst-case group accuracy of all ensemble methods also increases. Additionally, AGRE-KD  outperforms or matches the performance of the deep ensembling model, where the last layer of each model in the ensemble is directly retrained. These findings highlight the effectiveness of using gradients to steer model training towards specific goals, such as bias mitigation.
@@
Figure 3: Results on the proportion of debiased teachers in the ensemble. We trained the student model using an ensemble of 5 teachers with different debiased teacher ratios within the ensemble ($\{\frac{1}{5}$, $\frac{2}{5}$, …, $\frac{5}{5}\}$). AGRE-KD  effectively upweights and favors the least biased teachers in the ensemble, while other ensemble methods rely on them and decrease the WGA. AGRE-KD maintains significantly higher WGA, despite having only a single debiased model in the ensemble..
6 Conclusion
@@
In this paper, we studied bias in ensemble knowledge distillation (KD) and demonstrated that, unlike deep ensemble models that reduce bias, traditional ensemble KD methods can amplify it. We proposed AGRE-KD, an adaptive gradient-based weighting method that improves group robustness in ensemble KD by guiding the student model to learn core features and boosting worst-group accuracy (WGA). Our experiments across several benchmarks demonstrated the effectiveness of our approach in distilling knowledge with reduced spurious correlations. While our results highlight AGRE-KD’s advantages over existing methods, several questions remain. First, this study focused on unsupervised KD using teachers’ logits alone, without access to group or class labels. Second, WGA improvements are less pronounced when all teachers in the ensemble are biased, suggesting an opportunity to exploit class labels in this setting to further boost WGA, e.g., by considering the teachers’ misclassifications. Finally, additional evaluations on more complex datasets, such as language tasks, are needed to validate the approach across more diverse applications.
Acknowledgement
The authors thank the Digital Research Alliance of Canada for computing resources. SEK is supported by CIFAR and NSERC DG (2021-4086) and UA by NSERC DG (2022-04006).
Appendix A Algortihm
@@
4:         Compute knowledge distillation loss $\ell^{t}_{i}(\theta)$ for sample $i$ between $S$ and $T_{t}$triangleright (Eq. 1).
5:         Compute biased model distillation loss $\ell^{b}_{i}(\theta)$ for sample $i$ between $S$ and $T_{b}$triangleright (Eq. 1).
6:         Compute gradient alignment $G_{i}^{t}=\langle\nabla\ell^{t}_{i}(\theta),\nabla\ell^{b}_{i}(\theta)\rangle$triangleright Dot product of normalized gradient vectors. Table 4 shows the importance of ignoring the magnitude of the gradients.
@@ Appendix B Hyperparameters
We train all models using standard hyperparameters from previous work (LaBonte et al.,, 2024; Kirichenko et al.,, 2022; Izmailov et al.,, 2022) and keep their value fixed across experiments. For the vision tasks, we used an initial learning rate of $1\times 10^{-3}$ with a cosine learning rate scheduler; we used a batch size of 32 and 100 for the Waterbirds and the CelebA datasets, respectively. For the CiviComments dataset, we use an initial learning rate of $1\times 10^{-5}$ with a linear learning rate scheduler, a batch size 16, and train for ten epochs. We keep all hyperparameters fixed to train the teacher and student models. For the optimizer, we used AdamW (Loshchilov et al.,, 2019) and SGD for the language and vision datasets, respectively, with a weight decay of $1\times 10^{-4}$. Our implementation uses PyTorch (Paszke et al.,, 2017, 2019), Torch Lightning (Falcon and team,, 2019), and Milkshake (LaBonte,, 2023).
Appendix C Supplemental experiments

### v2

@@ 5.2.2 Results on real-world datasets
Teacher _ $85.7_{\pm 1.80}$ $65.6_{\pm 5.75}$ $95.4_{\pm 0.07}$ $37.5_{\pm 2.27}$ $90.0_{\pm 0.25}$ $75.9_{\pm 1.15}$
_ $94.2_{\pm 0.53}$ $90.9_{\pm 1.00}$ $93.7_{\pm 2.44}$ $90.1_{\pm 1.68}$ $85.9_{\pm 0.49}$ $77.9_{\pm 0.49}$
Deep Ensemble _ $84.4_{\pm 0.00}$ $59.3_{\pm 0.44}$ $95.6_{\pm 0.02}$ $37.2_{\pm 0.00}$ $90.6_{\pm 0.02}$ $76.1_{\pm 0.10}$
_ $93.5_{\pm 0.17}$ $90.0_{\pm 0.56}$ $92.4_{\pm 0.18}$ $90.3_{\pm 0.55}$ $85.9_{\pm 0.17}$ $76.6_{\pm 0.09}$
One-hot $\underline{83.8}_{\pm 0.97}$ $54.1_{\pm 2.21}$ $\underline{95.5}_{\pm 0.07}$ $32.3_{\pm 2.66}$ $90.1_{\pm 0.21}$ $\underline{75.5}_{\pm 0.84}$
$91.4_{\pm{1.31}}$ $86.7_{\pm 1.85}$ $92.3_{\pm 0.39}$ $88.9_{\pm 2.90}$ $85.7_{\pm 0.47}$ $\textbf{76.5}_{\pm 1.09}$
Random $80.0_{\pm 0.37}$ $39.6_{\pm 3.63}$ $95.2_{\pm 0.03}$ $27.2_{\pm 0.00}$ $90.8_{\pm{0.21}}$ $75.2_{\pm 0.95}$
@@
AGRE-KD (Ours) $82.2_{\pm 1.37}$ $\underline{55.0}_{\pm 5.47}$ $95.4_{\pm 0.04}$ $\underline{37.6}_{\pm 0.78}$ $89.3_{\pm 3.65}$ $74.7_{\pm 3.00}$
$\textbf{91.3}_{\pm 0.49}$ $\textbf{87.9}_{\pm 1.23}$ $91.7_{\pm 0.20}$ $\textbf{91.9}_{\pm 0.71}$ $90.2_{\pm 0.49}$ $75.9_{\pm 1.75}$
@@
AGRE-KD $84.9_{\pm 1.40}$ $\underline{66.3}_{\pm 4.76}$ $94.5_{\pm 3.22}$ $\underline{39.2}_{\pm 0.78}$
[…]
Table 3: Results on self-distillation. The student and the teacher models have the same network architecture (ResNet-50). The student’s worst group test accuracy increases when we distill to a network with higher capacity.
5.2.3 Impact of the model capacity.
@@
In this experiment, we study whether the student’s network capacity is a source of the implication of spurious feature learning in ensemble KD. The capacity here represents the number of trainable parameters of the model. For example, our ResNet-50 teacher models have around 25.6M parameters, substantially larger than ResNet-18 student models with approximately 11M parameters. We perform the same experiments as previously, but using self-distillation, i.e., we use the same architecture for the student and teacher models (e.g., ResNet-50 on image tasks). Results in Table 3 show that ensemble KD methods with self-distillation significantly improve the worst-group test accuracy. The gap between the teacher models and students is reduced compared to KD settings in Table 2, where the students’ models have smaller capacity. The results indicate that the students’ higher capacity can help the network learn more core features and reduce the influence of spurious features in the last layer. On the other hand, AGRE-KD outperforms other baselines, showing that our adaptive weighting scheme effectively guides the student models to focus on minimizing worst-case group error during training. We further illustrate the effectiveness of our weighting scheme in the next experiment by adjusting the number of debiased teachers in the ensemble.
5.2.4 Effect of the number of debiased teachers.
@@
Figure 3: Results on the proportion of debiased teachers in the ensemble. We trained the student model using an ensemble of 5 teachers with different ratios of debiased teachers within the ensemble ({$0.2$, $0.4$, $0.6$, $0.8$, 11}). AGRE-KD effectively upweights and favors the least biased teachers in the ensemble, while other ensemble methods rely more on biased teachers’ output and decrease the WGA. AGRE-KD maintains significantly higher WGA, despite having only a single debiased model in the ensemble.
In this experiment, we study how an ensemble with different proportions of biased and debiased models impacts the WGA of the student model. We use the same training process as previously and consider an ensemble of five teachers with different proportions of debiased teachers (i.e., {$0.2$, $0.4$, $0.6$, $0.8$, 11}); we report the average and WGA in Figure 3 for the Waterbirds and CelebA datasets. The results show that students’ WGA increases as the proportion of debiased models in the ensemble increases. When the ensemble contains a single debiased model, the aggregation process of other ensemble KD methods relies more on the majority of biased teachers, leading the student model to capture spurious correlation. On the other hand, our method can implicitly identify and adaptively rely on the knowledge of debiased teachers while reducing reliance on biased teachers. This is reflected by the significantly higher WGA in settings where the ensemble contains a single debiased teacher. Additionally, AGRE-KD outperforms or matches the performance of the deep ensembling model, where the last layer of each model in the ensemble is directly retrained.
5.2.5 Ablation on teacher architecture heterogeneity
Models Architecture KD Debiased Teachers ERM Teachers
Avg WGA Avg WGA
Teacher ViT $96.1_{\pm 0.30}$ $90.1_{\pm 1.50}$ $92.7_{\pm 0.00}$ $73.8_{\pm 1.50}$
ResNet-50 $94.3_{\pm 0.8}$ $9.5_{\pm 2.5} 88$$8.7_{\pm 0.09}$ $68.0_{\pm 2.20}$
ConvNeXtV2 $92.4_{\pm 0.6}$ $85.7_{\pm 1.5}$ $86.4_{\pm 0.41}$ $62.0_{\pm 0.55}$
RANDOM ResNet-18 $\textbf{94.3}_{\pm 0.18}$ $0.6_{\pm 2.30} 88$$8.9_{\pm 0.1}$ $57.7_{\pm 3.21}$
AVER $94.2_{\pm 0.21}$ $80.7_{\pm 1.10}$ $84.2_{\pm 0.4}$ $41.7_{\pm 2.51}$
AE-KD $93.5_{\pm 0.65}$ $84.9_{\pm 2.81}$ $89.1_{\pm 1.4}$ $57.9_{\pm 1.90}$
AGRE-KD $93.5_{\pm 0.63}$ $\textbf{87.9}_{\pm 1.25}$ $\textbf{89.5}_{\pm 0.7}$ $\textbf{61.6}_{\pm 2.10}$
Table 4: Results with heterogeneous teacher ensemble on the Waterbirds dataset. We train teachers with different architectures (ViT (base), ResNet-50, and ConvNeXtV2 (base)) and perform knowledge distillation to a ResNet-18 student model.
Experiments in Section 5.2.3 revealed that the capacity of the student model plays an important role in bias amplification of ensemble KD. Prior work has demonstrated that increasing the diversity of the ensembles results in better generalization performance. We enforced diversity by initializing the weights with independent random seeds (Allen-Zhu & Li, 2020). We now explore the heterogeneity of teacher models by building ensembles with diverse network architectures and analyzing the impact on the WGA accuracy of the student models. We train teacher models using three different network architectures, i.e., ViT (Dosovitskiy et al., 2020), ConvNeXtV2 (Woo et al., 2023), and ResNet-50 (He et al., 2016). Using the same experimental setup as previously, we built an ensemble of nine models with an equal proportion of each model architecture, i.e., three models with each architecture, all trained using different random initialization. This ensures diversity between and within teacher architectures.
The results in Table 4 show, for the Waterbirds dataset, the average and WGA accuracy when the heterogeneous ensemble contains ERM-trained or debiased teacher models. We observe that even under a more heterogeneous teacher ensemble, student models trained with classical ensemble KD methods can match the teacher’s average accuracy but incur a significant drop in WGA. Compared to student models trained with homogeneous ensembles of ResNet-50 models (Table 2), we observe that using a more diverse teacher’s backbone marginally improves performance across baselines, yet the drop in WGA remains important across ensemble KD methods. This highlights that the bias amplification in ensemble KD does not necessarily depend on ensemble diversity, further supporting our hypothesis of student capacity as a source of the disparity. On the other hand, the orchestration process in AGRE-KD shows better performance as its distillation process encourages better WGA. Despite the teachers’ architecture being heterogeneous, we use a ResNet-50 biased model to compute teachers’ weights in AGRE-KD during the distillation process, demonstrating the robustness of the proposed approach to the choice of teacher model architecture. The results on the CelebA dataset show the same trends, which we provide in Table 8 in the Appendix.
6 Limitation and conclusion
@@
In this paper, we studied bias in ensemble knowledge distillation (KD) and demonstrated that, unlike deep ensemble models that reduce bias, traditional ensemble KD methods can amplify it. We proposed AGRE-KD, an adaptive gradient-based weighting method that improves group robustness in ensemble KD by guiding the student model to learn core features and boosting worst-group accuracy. Our experiments across several benchmarks demonstrated the effectiveness of our approach in distilling knowledge with reduced spurious correlations. While our results highlight AGRE-KD ’s advantages over existing methods, several limitations and open questions remain. Specifically, the study focused on unsupervised KD using the teachers’ logits alone, without access to group or class labels. Although our primary interest was in examining the ability of an ensemble of teachers to distill more robust knowledge, exploring scenarios where class labels are available could enhance the performance of the student further. Moreover, we observed that the WGA improvement in the proposed method is less pronounced when all teachers in the ensemble are biased. This indicates an opportunity to leverage class labels in this context to further enhance group robustness, e.g., by considering the teachers’ misclassifications. Finally, additional evaluation using more complex datasets, such as health datasets, is necessary to validate the approach across a wider range of applications.
Appendix A AGRE-KD: Algorithm and method overview
In Algorithm 1, we provide a high-level description of the AGRE-KD  methodology for training and validation of group-robust distilled models. We also provide an overview of our method in Figure 2.
@@ Appendix A AGRE-KD: Algorithm and method overview
4:   Compute knowledge distillation loss $\ell^{t}_{i}(\theta)$ for sample $i$ between $S$ and $T_{t}$ $\triangleright$ Equation 1.
5:   Compute biased model distillation loss $\ell^{b}_{i}(\theta)$ for sample $i$ between $S$ and $T_{b}$ $\triangleright$ Equation  1.
6:   Compute gradient alignment $G_{i}^{t}=\langle\nabla\ell^{t}_{i}(\theta),\nabla\ell^{b}_{i}(\theta)\rangle$ $\triangleright$ Dot product of normalized gradient vectors. Table 6 shows the importance of ignoring the magnitude of the gradients.
@@ Appendix B Hyperparameters
We train all models using standard hyperparameters from previous work (LaBonte et al., 2024; Kirichenko et al., 2022; Izmailov et al., 2022) and keep their value fixed across experiments. For the vision tasks, we used an initial learning rate of $1\times 10^{-3}$ with a cosine learning rate scheduler; we used a batch size of 32 and 100 for the Waterbirds and the CelebA datasets, respectively. For the CivilComments dataset, we use an initial learning rate of $1\times 10^{-5}$ with a linear learning rate scheduler, a batch size 16, and train for ten epochs. We keep all hyperparameters fixed to train the teacher and student models. For the optimizer, we used AdamW (Loshchilov et al., 2019) and SGD for the language and vision datasets, respectively, with a weight decay of $1\times 10^{-4}$. Our implementation uses PyTorch (Paszke et al., 2017; 2019), Torch Lightning (Falcon & team, 2019), and Milkshake (LaBonte, 2023).
Appendix C Ablation studies
C.1 Ablation on $\alpha$
Equations 3 and 7 define the student training objective as a weighted combination of the cross-entropy loss with ground-truth labels and the ensemble knowledge distillation loss. The hyperparameter $\alpha\in[0,1]$ controls this weighting: smaller values place more emphasis on the ground-truth loss, while larger values prioritize the distillation signal. Since our primary focus is to study how bias propagates through ensemble KD, all main paper experiments fixed $\alpha=1.0$, thereby relying entirely on distillation.
To investigate the impact of $\alpha$, we performed an ablation with $\alpha\in\{0.1,0.3,0.5,0.7,0.9,1.0\}$ on the Waterbirds and CelebA datasets. We used ensembles of five ResNet-50 DFR teachers and a ResNet-18 student, and report results averaged over three random seeds. Figure 4 presents the trends for both average accuracy and worst-group accuracy (WGA).
We observe that WGA increases nearly monotonically with $\alpha$, indicating that when the KD loss dominates, students inherit more of the debiasing effect from the DFR teachers. For example, in Waterbirds with AGRE-KD, WGA rises from $62.5\%$ at $\alpha=0.1$ to $84.6\%$ at $\alpha=1.0$. The best-performing baseline (Average KD) improves from $62.3\%$ to $81.3\%$ across the same range, yet still falls short of the One-hot model trained directly with DFR ($83.9\%$ WGA).
These results reinforce two key points: (i) ensemble KD benefits from stronger reliance on distillation when teachers are debiased, and (ii) without dedicated mechanisms such as AGRE-KD, ensemble KD can lag behind simpler approaches like direct DFR training. This highlights the importance of explicitly addressing bias amplification in ensemble distillation.
Figure 4: Effect of the parameter $\alpha$ (Eq. 3) on the worst-group accuracy of DFR teachers.
C.2 Ablation on temperature parameter
Our experiments using the DFR teacher models revealed that reweighting the feature by retraining the last classification layer provides a "dark knowledge" (soft target) to the student model that is more robust to spurious correlation. Since DFR teachers share the same feature representation as their ERM-trained counterparts but differ only in their final classification layer, the enhanced robustness of the student when using DFR teachers indicates that the probability distribution over all classes predicted by the teacher significantly influences the student’s dependence on spurious correlations. Since the temperature hyperparameter (Equation 1) softens the probability distribution of the teacher to better expose dark knowledge. We perform an ablation study by varying the temperature parameter ($\tau\in\{1,2,4,...,10\}$) and analyze its influence on the spurious correlation captured by the student model.
Figure 5 shows the WGA averaged across three independent random seeds in WaterBirds and CelebA datasets. We considered ensembles of five ResNet-50 DFR teachers and a ResNet-18 student. The results show that WGA increases with temperature and then remains steady or decreases after higher values of temperature ($\tau\geq 10$ and $\tau\geq 8$ for the Waterbirds and CelebA datasets, respectively). The smallest WGA is achieved when $\tau=1$, which means the training uses a regular cross-entropy loss; it increases when we soften the probability distribution ($\tau>1$) and eventually decreases when the distribution gets almost uniform ($\tau\gg 1$). These findings suggest that the dark knowledge provided by the softened probability distribution of the teacher models reduces students’ reliance on spurious correlation during training. These results align with recent work by Mohammadshahi & Ioannou (2024) studying the benefit of increased temperature for fairness in knowledge distillation.
Figure 5: Effect of the temperature parameter on the worst-group accuracy.
Appendix D Supplemental results